OpenStudy (anonymous):

Please check my first 4 answers and help with the last question?!

OpenStudy (anonymous):

Question. What is the mole fraction of K in 55.3 g KBr dissolved in 350 g of water (H2O)? Atomic masses: K= 39.0 g/mol; Br= 79.9 g/mol H=1.0 g/mol O=16.0 g/mol a) What is the solute’s molar mass? b) Convert grams of solute to moles of solute using molar mass as a conversion factor. c) Determine the solvent’s molar mass. d) Convert grams of solvent to moles of solvent using molar mass as a conversion factor. e) What is the mole fraction of K? Answer. a) 39.0 g/mol + 79.9 g/mol = 118.9 g/mol b) 55.3 g / 118.9 g/mol = 0.465 mol c) (1.0*2) + 16.0 g/mol = 18.0 g/mol d) 350 g / 18.0 g/mol = 19.4 mol e) ?

OpenStudy (anonymous):

@rock_mit182

OpenStudy (rock_mit182):

let me remember something mole fraction (x) = a mol/ (a+b) mol ?

OpenStudy (rock_mit182):

seriously i don't know by memory the formulas